OpenStudy (anonymous):

what is the equation of a line that passes through (0, -2) and (2, 3)

OpenStudy (ranga):

Find the slope m from the two points: m = (y2 - y1) / (x2 - x1) Then the equation of the line is: (y - y1) = m(x - x1)